40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3181     Entry time: Thu Jul 8 17:29:20 2010
Author: Katharine, Sharmila 
Type: Update 
Category: elog 
Subject:  


Last night, we successfully connected and powered our circuit, which allowed us to test whether our OSEMs were working.  Previously, we had been unable to accomplish this because (1) we weren't driving it sufficiently high voltage, and  (2) we didn't check that the colored leads on our circuit actually corresponded to the colored ports on the power supply (they were all switched, which we are in the process of rectifying), so our circuit was improperly connected to the supply .  Unfortunately, we didn't learn this until after nearly cooking our circuit, but luckily there appears to have been no permanent damage .

Our circuit specs suggested powering it with a voltage difference of 48V, so we needed to run our circuit at a difference of at least 36-40 V.  Since our power supply only supplied a difference of up to 30V in each terminal, we combined them in order to produce a voltage of up to double that.  We decided to power our circuit with a voltage difference of 40V (+/- 20V referenced to true ground).  The current at the terminals were 0.06 and 0.13 A. 

To test our circuit, we used a multimeter to check the supplied voltage at different test points, to confirm that an appropriate input bias was given to various circuit elements.  We identified the direction of LED bias on our OSEM, and connected it to our circuit. We were extremely gratified when we looked through the IR viewer and saw that, in fact, the LED in the OSEM was glowing happily .

P7070240.JPG P7070242.JPG

We hooked up two oscilloscopes and measured the current through the coil, and also through the LED and photodiode in the OSEM.  We observed a change in the photodiode signal when we blocked the LED light, which was expected.  The signal at the PD and the LED were both sinusoidal waves around ~3 kHz.

P7070255.JPG 

P7070257.JPG

 

We then went back to our levitation setup, and crudely tried to levitate a magnet with attached flag by using our hands and adjusting the gain (though we also could have been watching the PD current).  The first flag we tried was a soldering tip; we couldn't levitate this but achieved an interesting sort of baby-step "levitation" (levitation .15) which allowed us to balance the conical flag on its tip on top of the OSEM (stable to small disturbances).  After learning that conical flags are a poor idea, we switched our flag to a smaller-radius cylindrical magnet.  We were much closer to levitating this magnet, but were unable to conclusively levitate it .

 P7070249.JPG
Current plan:

Adjust the preset resistors to stabilize feedback

Check LED drive circuit.

Finish calculating the transfer function, and hook up the circuit to the spectrum analyzer to measure it as well.

Observe the signal from the photocurrent as disturbances block the LED light.

Play with the gain of the feedback to see how it affects levitation.

 

Attachment 1: P7070254.JPG  4.506 MB  | Hide | Hide all
P7070254.JPG
ELOG V3.1.3-